OpenStudy (anonymous):

Solve 2x + 5 = 27 @saue7414

OpenStudy (anonymous):

x=11

OpenStudy (anonymous):

Simplifying 2x + 5 = 27 Reorder the terms: 5 + 2x = 27 Solving 5 + 2x = 27 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-5' to each side of the equation. 5 + -5 + 2x = 27 + -5 Combine like terms: 5 + -5 = 0 0 + 2x = 27 + -5 2x = 27 + -5 Combine like terms: 27 + -5 = 22 2x = 22 Divide each side by '2'. x = 11 Simplifying x = 11

OpenStudy (anonymous):